# Stability of quasicrystalline ultracold fermions to dipolar interactions

### v1

@@ Abstract
Quasiperiodic potentials can be used to interpolate between localization and delocalization in one dimension. However, little is known about the stability of quasicrystalline phases to long-range interactions. In this work, we study repulsive ultracold dipolar fermions in a quasiperiodic optical lattice to characterize the behavior of interacting quasicrystals. We simulate the full time evolution of the typical experimental protocols used to probe quasicrystalline order and localization properties. We extract experimentally measurable dynamical observables and correlation functions to characterize the three phases observed in the noninteracting setting: localized, intermediate, and extended. We then study the stability of such phases to repulsive dipolar interactions. We find that dipolar interactions can completely alter the shape of the phase diagram by stabilizing the intermediate phase, mostly at the expense of the extended phase. Moreover, in the strongly interacting regime, a resonance-like behavior characterized by density oscillations appears. Remarkably, strong dipolar repulsions can also localize particles even in the absence of quasiperiodicity if the primary lattice is sufficiently deep. Our work shows that dipolar interactions in a quasiperiodic potential can give rise to a complex, tuneable coexistence of localized and extended quantum states.
@@
Introduction — Since their revolutionary discovery in aluminum-manganese alloys Shechtman et al. (1984), quasicrystals have been the subject of intense theoretical and experimental research due to their many exotic properties. Quasicrystals exhibit a regular structure despite lacking a canonical periodicity Rabson et al. (1991); Goldman and Kelton (1993). While they lack translational symmetry, they can completely tesselate a two-dimensional surface in an aperiodic fashion because they do possess unexpected rotational symmetries (five-fold, eight-fold, etc.). Quasicrystals can be mathematically understood in terms of higher-dimensional periodic structures projected onto lower-dimensional slices, and have thus ushered in new paradigms for crystallography Mackay (1995); Yamamoto (1996); Steurer and Deloudi (2008, 2009); Wolny et al. (2018), band theory in solids Poon (1992); Kelton (1993); Rotenberg et al. (2004); Grimm and Schreiber (2003), and hidden dimensions Jagannathan (2021). Quasiperiodic structures can further emulate on-site disorder, which makes them ideal platforms to study many-body localization Iyer et al. (2013); Schreiber et al. (2015); Setiawan et al. (2017); Khemani et al. (2017); Zhang and Yao (2018); Macé et al. (2019); Varma and Žnidarič (2019); Abanin et al. (2019); Agrawal et al. (2022) and its interplay with nonergodicity Li et al. (2015, 2016); Hsu et al. (2018); Ghosh et al. (2020); Roy et al. (2023), integrability He et al. (2013); Žnidarič and Ljubotina (2018); Singh et al. (2021); Žnidarič (2021); Thomson and Schirò (2023), and fractality Thiem and Schreiber (2011); Kalugin and Katz (2014); dimensions of wave functions et al. (2016); Macé et al. (2017); Rai et al. (2019, 2020); Jagannathan (2021); Dai et al. (2023).
@@
Impressive advances in trapping and controlling ultracold atoms have made them an excellent platform where to realize and study quasiperiodicity with extreme precision Lye et al. (2007); Fallani et al. (2007); Singh et al. (2015); Schreiber et al. (2015); Macé et al. (2016); Bordia et al. (2017); Lüschen et al. (2017a, b); Lueschen et al. (2018); Corcovilos and Mittal (2019); Viebahn et al. (2019); Johnstone et al. (2019); Sbroscia et al. (2020); Gautier et al. (2021); Nakajima et al. (2021); Yu et al. (2023). In one dimension (1D), a quasiperiodic potential can be generated by superimposing a primary optical lattice with a detuning lattice at an incommensurate frequency. Upon loading particles in it, it is possible to quantum simulate and study the dynamics of 1D quasicrystals with high accuracy. A particularly interesting feature observed is the coexistence of localized and extended states in an intermediate phase separating the fully localized to the fully extended phase Lueschen et al. (2018). In the intermediate phase, the different states are separated by a critical energy level termed single-particle mobility edge (SPME) Abrahams et al. (1979); Lee and Ramakrishnan (1985); Boers et al. (2007); Ganeshan et al. (2015); Li et al. (2017); Lueschen et al. (2018). This coexistence manifests itself due to long-range tunneling terms of the continuum description of the 1D ultracold quantum simulator. In fact, the SPME is absent in the deep lattice limit, where an effective Aubry-André model is obtained Aubry and André (1980); Fallani et al. (2007); Roati et al. (2008); Modugno (2009). Furthermore, a rigorous SPME does not appear in 1D systems with Anderson-type (uncorrelated) disorder Lueschen et al. (2018).
[…]
@@
In this work, we study the role of DDI in 1D ultracold fermions loaded in a quasiperiodic potential. We simulate the exact experimental protocol Schreiber et al. (2015); Lueschen et al. (2018); Kohlert et al. (2019); An et al. (2021), whereby fermions are first loaded in a superlattice of twice the periodicity of the primary optical lattice, and then are let to time evolve into the quasiperiodic structure quenched at time zero. By measuring the imbalance between odd and even sites and the expansion of the particle density, it is possible to classify the system as being in a localized, extended, or intermediate phase. We study the stability of each phase against DDI by determining phase diagrams and calculating observables such as pair-correlation functions. We find that DDI stabilize the intermediate phase, mostly to the detriment of the extended phase. This stabilization is a result of a decreased correlation between far away particles that can be attributed to interference between the long-range tails of the DDI. At strong DDI, the same mechanism not only strengthens the imbalance in the localized phase, but also induce a new kind of localization in regimes where the noninteracting system would only host extended states. Surprisingly, we find that this localization persists also when the detuning lattice is zero, i.e. for a fully periodic optical lattice. Our study illustrates the potential for long-range interacting systems to induce new types of localization phenomena in conjunction with quasicrystalline structures.
@@
with $E_{r}$ the recoil energy of the primary lattice (unit of energy). For the potential parameters we choose values compatible with experimental realizations of this system Lueschen et al. (2018); sup . We also add hard wall boundaries Gaunt et al. (2013); Mazurenko et al. (2017); Gall et al. (2021); Navon et al. (2021) to obtain a finite system containing 64 sites in the primary lattice  [^1] While this spatial extent is smaller than in typical experiments, it is more than enough to capture the hallmarks of each phase.. The fermions interact with one another with the interaction Hamiltonian
@@
is a dipolar repulsion with interaction strength $W$ and regularization factor $\alpha=0.01$  [^2] This phenomenologically accounts for Pauli pressure., and $\bar{L}$ is the distance between two neighboring minima in the primary lattice (unit of length).
@@
with $\mathcal{N}_{c}=\int_{c}\mathrm{d}x\>\rho(x,t)$ and $c$ referring to the initially populated central region (e.g. for $N=4$, $c=[-4.5,3.5]$). These two observables in the long-time limit determine the many-body phase of the system. A finite imbalance $\mathcal{I}$ indicates the persistence of the initial superlattice charge density wave pattern and is a proxy for the presence of localized states. A finite (and growing) value of the edge density $\mathcal{D}$ reveals instead the presence of extended states. Therefore, in the localized phase $\mathcal{I}>0$ and $\mathcal{D}=0$, while in the extended phase $\mathcal{I}=0$ and $\mathcal{D}>0$. Localized and extended states can also coexist in an intermediate phase, where both values of $\mathcal{I}$ and $\mathcal{D}$ are nonzero and a SPME is present. While the SPME should occur at arbitrarily small nonzero $\mathcal{I}$ and $\mathcal{D}$, numerical simulations introduce small imprecisions. Thus small but finite thresholds $t_{\mathcal{I}}$ and $t_{\mathcal{D}}$ for the imbalance and the edge density have to be introduced. Throughout our analysis, we will empirically set them to be $t_{\mathcal{I}}=0.1$ and $t_{\mathcal{D}}=0.03$  [^3] We found empirically that these threshold values minimize numerical deviations from the experimentally observed phase diagram..
We also probe the correlations within each phase. For simplicity, we focus here on the behavior of the diagonal of the reduced two-body density matrix (2-RDM)
@@
where $\left|\Psi(t)\right>$ is the time-evolved many-body state of the system. This is equivalent to the pair-correlation function and describes the conditional probability of finding a fermion at position $x$, when another fermion is located at $x^{\prime}$, i.e. it quantifies how a fermions is surrounded by other fermions. Thus, it encodes the entire information about pairwise interactions among particles and is a proxy for the (unnormalized) pairwise correlations of the system.
@@
Methods — To simulate the full many-body interacting system, we use the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles Streltsov et al. (2006, 2007); Alon et al. (2007, 2008), coded in the MCTDH-X software Lode (2016); Fasshauer and Lode (2016); Lin et al. (2020a); Lode et al. (2020, 2024). With it, we solve the many-body Schrödinger equation directly for the continuum system consisting of kinetic energy, interparticle interactions, and optical lattice. MCTDH-X has been widely used to study ground-state and dynamical properties of long-range interacting systems Fischer et al. (2015); Molignini et al. (2018); Chatterjee and Lode (2018); Lin et al. (2019); Chatterjee et al. (2019, 2020); Lin et al. (2020b, 2021); Molignini et al. (2022); Hughes et al. (2023); Bilinskaya et al. (2024); Molignini and Chakrabarti (2024a, b). MCTDH-X relies on a time-dependent variational optimization procedure in which the many-body wavefunction is decomposed into an adaptive basis set of $M$ time-dependent single-particle functions called orbitals. For strongly interacting systems, a larger number of orbitals might be required to capture many-body correlation effects. We have verified that including orbitals beyond $M=N+4$ leads to a negligible population for the time scales probed by our dynamics sup .
@@
Results for noninteracting fermions — We begin by describing the dynamics for noninteracting particles. Fig. 1(a) shows the phase diagram for $N=4$ fermions with $M=8$ orbitals, constructed from imbalance and expansion as the depth of the primary and detuning lattices are varied. We use experimentally achievable parameter ranges, i.e. $V_{p}\in[3E_{r},8E_{r}]$ and $V_{d}\in[0,1.2E_{r}]$. The imbalance is calculated by averaging over the last 50 time steps in the time evolution. The edge density is calculated at a target time $\bar{t}$, well before the particles can reach the boundaries of the system. Typical target times are $\bar{t}=0.45$ ms to $\bar{t}=1.78$ ms  [^4] This measurement timescale is slightly different than the experimental one, where expansion measurements are calculated at longer times. This is necessary to avoid simulating very large spatial grids that are computationally unattainable. Nevertheless, we have verified that the value of the target time $\bar{t}$ only slightly impacts the precise location of the phase boundary between intermediate and localized phase, and that the qualitative behavior is the same for different values of $\bar{t}$.. From Fig. 1(a), we can clearly distinguish the three expected phases, which have the same structure observed in prior studies Li et al. (2017); Lueschen et al. (2018). At low values of $V_{d}$, the system is always in the extended phase (purple region). Increasing the value of $V_{d}$ at low to moderate values of $V_{p}$ pushes the system into the intermediate phase (magenta region). Further increasing $V_{d}$ or $V_{p}$ eventually localizes all the states (yellow region).
@@
Figure 1: Behavior of the noninteracting system: (a) Phase diagram for $N=4$ fermions in a 1D quasiperiodic potential. For the localized phase, the time-averaged imbalance $\mathcal{I}\right>$ is plotted in a continuous color scheme. (b)-(d) 2-RDM at $t=0.89$ ms for (a) localized phase ($V_{p}=8.0E_{r}$, $V_{d}=1.2E_{r}$), (b) intermediate phase ($V_{p}=6.0E_{r}$, $V_{d}=0.4E_{r}$), and (c) extended phase ($V_{p}=3.0E_{r}$, $V_{d}=0E_{r}$).
@@
The three phases exhibited by the quasicrystal system can also be distinguished by the dynamics of their many-body correlations encoded in the 2-RDM, shown for selected times in Fig. 1(b)-(d) (see supplementary materials for more time snapshots sup ). The 2-RDM in the localized phase [Fig. 1(b)] does not change at all with time and merely shows the expected exchange hole at $x=x^{\prime}$ with strong many-body correlation between the sites of the initial superlattice. On the contrary, the extended phase [Fig. 1(d)] exhibits a rapid spread of correlations at very short times already. Between the two extremes of localization and expansion lies the intermediate phase [Fig. 1(c)]. The correlation pattern here reveals an almost immediate exchange correlation between neighboring sites within the populated area of the initial superlattice, and a small spread to sites just outside of it. However, the intensity of the 2-RDM is not uniform: correlations between sites in the initial superlattice structure still dominate throughout the time evolution even for longer times sup . These results indicate that the intermediate phase is characterized by two types of correlations showcasing the coexistence of localized and extended states. Localized states mediate static, short-range correlations between the initially populated sites, while extended states mediate long-range correlations spanning many sites in the quasiperiodic lattice.
@@
Fig. 2 shows how the noninteracting phase diagram changes when interactions are progressively increased [panels (a) to (c)]. At small DDI, the intermediate phase only grows slightly while the imbalance in the localized phase decreases. As the DDI become stronger, though, the phase diagram changes drastically and in a non-monotonic fashion. For $W=0.2E_{r}$, the intermediate phase grows sizeably at the expense of most of the localized phase and part of the extended phase. Thus, DDI stabilize the coexistence of extended and localized states in the intermediate phase. At even stronger DDI, $W=2.0E_{r}$, the intermediate phase dominates the probed parameter region. Note that in Fig. 2(c) $V_{p}$ is plotted to larger values up to $12E_{r}$. The extended phase is reduced to a tiny pool around $V_{p}=3E_{r}$, $V_{d}=0E_{r}$. Most strikingly, however, a new localized region emerges at low values of $V_{d}$, a region that is part of the extended phase in the noninteracting case. The localized phases at low and high $V_{d}$ are divided by a strip of intermediate phase which persists for very large values of $V_{p}$  [^5] We verified that the intermediate phase lobe persists up to values as large as $V_{p}=16.0E_{r}$ – not shown in these plots., reminiscent of the Arnold tongues found in parametrically driven systems McLachlan (1951); Magnus and Winkler (1966); Molignini et al. (2018). In this lobe, the particles undergo density oscillations around their initial pinned configuration, indicating a resonance between DDI and quasiperiodic potential [see Fig. 2(f)]. While both regions above and below the intermediate phase lobe show strong localization, their behavior is different. Above the lobe (high $V_{d}$, Fig. 2(d)), the particles are frozen in their initial configuration. Below the lobe (low $V_{d}$, Fig. 2(e)), the particle density periodically leaks back and forth from the initial boundaries while retaining a strong localization inside them, indicating a dynamically self-bound state.
@@
A more systematic calculation of the density dynamics, shown in Fig. 3, allows us to better discuss the behavior in each phase. In the extended phase [Fig. 3(a), (d), (g)], the DDI have initially little effect on the density, which maintains its linear spread in time consistent with the Lieb-Robinson bounds Tran et al. (2021). At strong interactions, however, the long-range repulsive tail in the DDI hinders the spread of correlations Bera et al. (2019), resulting in both a delayed expansion and a more chaotic density spread [Fig. 3(g)]. In the localized phase [Fig. 3(b), (e), (h))] tiny oscillations can be seen at one edge of the density profile up to intermediate DDI. At strong DDI, however, this behavior is suppressed. This stabilization of localization is also observed in the intermediate phase [Fig. 3(c), (f), (i)]. At low DDI, the density in the intermediate phase maintains a strong imbalance towards the initial superlattice configuration (due to localized states) while expanding towards outer lattice sites (due to extended states). However, stronger DDI increase the imbalance while simultaneously reducing the expansion, again consistent with a less effective spread of correlations in dipolar systems due to the interference between repulsive DDI tails. This is also an indication that DDI can be used to control the value of the SPME similarly to what was shown for contact interactions Wang et al. (2022).
@@
The increase in localization due to DDI interference is also observed in the 2-RDM, shown in Fig. 4. The behavior for weak DDI [$W=0.02E_{r}$, Fig. 4(a)-(c)] is similar to what already observed in the noninteracting case. However, when DDI are increased [$W=0.2E_{r}$ in Fig. 4(d)-(f) and $W=2.0E_{r}$ in Fig. 4(g)-(i)], the correlation spread becomes progressively stiffer for all phases. This leads to a complete freeze of correlations in the localized phase [Fig. 4(g)], while in the intermediate phase correlations are able to spread only at the right boundary  [^6] We attribute this unidirectional spread to pinning by the given choice of the quasiperiodic potential. Thus, the location of the pinning might change with a different choice of the wave vector ratio and detuning phase.. Even for the extended phase, the pinning effect is visible, with correlations spreading more easily towards the right than the left. This observation agrees with the uneven density spread observed in Fig. 3(g).

### v2

@@ Abstract
Quasiperiodic potentials can be used to interpolate between localization and delocalization in one dimension. With the rise of optical platforms engineering dipolar interactions, a key question is the stability of quasicrystalline phases under these long-range interactions. In this work, we study repulsive ultracold dipolar fermions in a quasiperiodic optical lattice to characterize the behavior of interacting quasicrystals. We simulate the full time evolution of the typical experimental protocols used to probe quasicrystalline order and localization properties. We extract experimentally measurable dynamical observables and correlation functions to characterize the three phases observed in the noninteracting setting: localized, intermediate, and extended. We then study the stability of such phases to repulsive dipolar interactions. We find that dipolar interactions can completely alter the shape of the phase diagram by stabilizing the intermediate phase, mostly at the expense of the extended phase. Moreover, in the strongly interacting regime, a resonance-like behavior characterized by density oscillations appears. Remarkably, strong dipolar repulsions can also localize particles even in the absence of quasiperiodicity if the primary lattice is sufficiently deep. Our work shows that dipolar interactions in a quasiperiodic potential can give rise to a complex, tuneable coexistence of localized and extended quantum states.
@@
Introduction — Since their revolutionary discovery in aluminum-manganese alloys Shechtman et al. (1984), quasicrystals have been the subject of intense theoretical and experimental research due to their many exotic properties. Quasicrystals exhibit a regular structure despite lacking a canonical periodicity Rabson et al. (1991); Goldman and Kelton (1993). While they lack translational symmetry, they can completely tesselate a two-dimensional surface in an aperiodic fashion because they do possess unexpected rotational symmetries (five-fold, eight-fold, etc.). Quasicrystals can be mathematically understood in terms of higher-dimensional periodic structures projected onto lower-dimensional slices, and have thus ushered in new paradigms for crystallography Mackay (1995); Yamamoto (1996); Steurer and Deloudi (2008, 2009); Wolny et al. (2018), band theory in solids Poon (1992); Kelton (1993); Rotenberg et al. (2004); Grimm and Schreiber (2003), and hidden dimensions Jagannathan (2021). Moreover, quasiperiodic structures can emulate on-site disorder, which makes them ideal platforms to study many-body localization Iyer et al. (2013); Schreiber et al. (2015); Setiawan et al. (2017); Khemani et al. (2017); Zhang and Yao (2018); Macé et al. (2019); Varma and Žnidarič (2019); Abanin et al. (2019); Agrawal et al. (2022) and its interplay with nonergodicity Li et al. (2015, 2016); Hsu et al. (2018); Ghosh et al. (2020); Roy and Sharma (2021); Roy et al. (2023), integrability He et al. (2013); Žnidarič and Ljubotina (2018); Singh et al. (2021); Žnidarič (2021); Thomson and Schirò (2023), and fractality Thiem and Schreiber (2011); Kalugin and Katz (2014); dimensions of wave functions et al. (2016); Macé et al. (2017); Rai et al. (2019, 2020); Jagannathan (2021); Dai et al. (2023).
@@
In recent years, impressive advances in trapping and controlling ultracold atoms have made them an excellent platform where to realize and study quasiperiodicity with extreme precision Lye et al. (2007); Fallani et al. (2007); Singh et al. (2015); Schreiber et al. (2015); Macé et al. (2016); Bordia et al. (2017); Lüschen et al. (2017a, b); Lueschen et al. (2018); Corcovilos and Mittal (2019); Viebahn et al. (2019); Johnstone et al. (2019); Sbroscia et al. (2020); Gautier et al. (2021); Nakajima et al. (2021); Yu et al. (2023). In one dimension (1D), a quasiperiodic potential can be generated by superimposing a primary optical lattice with a detuning lattice at an incommensurate frequency. Upon loading particles in such a quasiperiodic potential, it is possible to quantum simulate and study the dynamics of 1D quasicrystals with high accuracy. A particularly interesting feature exhibited by these systems is the coexistence of localized and extended states in an intermediate phase separating the fully localized to the fully extended phase Lueschen et al. (2018). In the intermediate phase, the different states are separated by a critical energy level termed single-particle mobility edge (SPME) Abrahams et al. (1979); Lee and Ramakrishnan (1985); Boers et al. (2007); Ganeshan et al. (2015); Li et al. (2017); Lueschen et al. (2018). This coexistence manifests itself due to long-range tunneling terms that appear in the continuum description of the 1D ultracold quantum simulator. In fact, the SPME is absent in the deep lattice limit, where an effective Aubry-André model with self-duality is obtained Aubry and André (1980); Fallani et al. (2007); Roati et al. (2008); Modugno (2009). Furthermore, a rigorous SPME does not appear in 1D systems with Anderson-type (uncorrelated) disorder Lueschen et al. (2018).
While the properties of the SPME and intermediate phase have been widely explored in noninteracting systems, less attention has been given to potential interplays with interactions, in particular long-range ones that can arise in the new generation of dipolar quantum simulators Micheli et al. (2006); Lahaye et al. (2009); Baranov et al. (2012); Gross and Bloch (2017). In fact, nowadays numerous atomic and molecular species exhibiting dipole-dipole interactions (DDI) are available in ultracold labs, such as dysprosium 161Dy Lu et al. (2012), erbium 167Er Aikawa et al. (2014); Baier et al. (2018), chromium53Cr Naylor et al. (2015), potassium-rubidium 40K87Rb Ni et al. (2008), sodium-lithium 23Na6Li Rvachov et al. (2017), and sodium-potassium 23Na40Park et al. (2015); Duda et al. (2023). DDI, beside better incorporating the long-range nature of interactions in solid-state quasicrystalline materials, could potentially herald many new interesting phases of matter when interfaced with quasicrystalline long-range order Bai and Xue (2015); Pandey et al. (2020); Vu et al. (2022).
@@
In this work, we investigate the role of DDI in 1D ultracold fermions loaded in a quasiperiodic potential. We simulate the exact experimental protocol used in ultracold atomic setups Schreiber et al. (2015); Lueschen et al. (2018); Kohlert et al. (2019); An et al. (2021), whereby the fermions are first loaded in a superlattice of twice the periodicity of the primary optical lattice, and then are let to time evolve into the quasiperiodic structure quenched at time zero. By measuring the imbalance between odd and even sites and the expansion of the particle density, it is possible to classify the state of the system as being in a localized, extended, or intermediate phase. We study the stability of each phase with respect to DDI by determining phase diagrams and by calculating observables such as pair-correlation functions.
We find that DDI tend to stabilize the intermediate phase, mostly to the detriment of the extended phase. Examining the correlation functions, this stabilization results from a decreased correlation between far away particles that can be attributed to interference between the long-range tails of the DDI. At strong DDI, the same mechanism not only strengthens the imbalance in the localized phase, but also induce a new kind of localization in regimes where the noninteracting system would only host extended states. Surprisingly, we find that this localization persists also when the detuning lattice is zero, i.e. for a fully periodic optical lattice. By combining the effect of quasiperiodicity and DDI-induced bound clusters, we bring forth an explanation for the pathway from disorder-free localization, to delocalization, to eventual re-localization. Our study illustrates the potential for long-range interacting systems to induce new types of localization phenomena in conjunction with quasicrystalline structures.
@@
with $E_{r}$ the recoil energy of the primary lattice (unit of energy). For the potential parameters we choose values compatible with experimental realizations of this system Lueschen et al. (2018); sup . We also add hard wall boundaries Gaunt et al. (2013); Mazurenko et al. (2017); Gall et al. (2021); Navon et al. (2021) to obtain a finite system containing 64 sites in the primary lattice. While this spatial extent is smaller than in typical experiments, it is more than enough to capture the hallmarks of each phase.
The fermions interact with one another with the interaction Hamiltonian
@@
is a dipolar repulsion with interaction strength $W$ and regularization factor $\alpha=0.01$ (phenomenologically accounting for Pauli pressure), and $\bar{L}$ is the distance between two neighboring minima in the primary lattice (unit of length).
@@
with $\mathcal{N}_{c}=\int_{c}\mathrm{d}x\>\rho(x,t)$ and $c$ referring to the initially populated central region (e.g. for $N=4$, $c=[-4.5,3.5]$). These two observables in the long-time limit help us to determine the many-body phase of the system. A finite imbalance $\mathcal{I}$ indicates the persistence of the initial charge density wave pattern imposed by the superlattice and is a proxy for the presence of localized states. A finite (and growing) value of the edge density $\mathcal{D}$ reveals instead the presence of extended states. Therefore, in the localized phase $\mathcal{I}>0$ and $\mathcal{D}=0$, while in the extended phase $\mathcal{I}=0$ and $\mathcal{D}>0$. Localized and extended states can also coexist at different energies in an intermediate phase, where both values of $\mathcal{I}$ and $\mathcal{D}$ are nonzero and a SPME is present. While the SPME should occur at arbitrarily small nonzero $\mathcal{I}$ and $\mathcal{D}$, numerical simulations introduce small imprecisions. Thus small but finite thresholds $t_{\mathcal{I}}$ and $t_{\mathcal{D}}$ for the imbalance and the edge density have to be introduced. Throughout our analysis, we will empirically set them to be $t_{\mathcal{I}}=0.1$ and $t_{\mathcal{D}}=0.03$  [^1] We found empirically that these threshold values minimize numerical deviations from the experimentally observed phase diagram..
Besides exploring how interactions impact the phase diagram and the boundaries between different phases, we also probe the correlations within each phase. For simplicity, we focus here on the behavior of the diagonal of the reduced two-body density matrix (2-RDM), defined as
@@
where $\left|\Psi(t)\right>$ is the time-evolved many-body state of the system. The diagonal of the 2-RDM is equivalent to the pair-correlation function and describes the conditional probability of finding a fermion at position $x$, when another fermion is located at $x^{\prime}$, i.e. it quantifies how a fermions is surrounded by other fermions. Thus, it encodes the entire information about pairwise interactions among particles and is a proxy for the (unnormalized) pairwise correlations of the system.
@@
Methods — To simulate the full many-body interacting system, we employ the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles Streltsov et al. (2006, 2007); Alon et al. (2007, 2008), implemented in the MCTDH-X software Lode (2016); Fasshauer and Lode (2016); Lin et al. (2020a); Lode et al. (2020, 2024). With it, we solve the many-body Schrödinger equation directly for the continuum system consisting of kinetic energy, interparticle interactions, and optical lattice. MCTDH-X has been widely used to study ground-state and dynamical properties of long-range interacting systems Fischer et al. (2015); Molignini et al. (2018); Chatterjee and Lode (2018); Lin et al. (2019); Chatterjee et al. (2019, 2020); Lin et al. (2020b, 2021); Molignini et al. (2022); Hughes et al. (2023); Bilinskaya et al. (2024); Molignini and Chakrabarti (2024a, b). MCTDH-X relies on a time-dependent variational optimization procedure in which the many-body wavefunction is decomposed into an adaptive basis set of $M$ time-dependent single-particle functions called orbitals. Due to the Pauli exclusion principle, to correctly describe a set of $N$ fermions, we require $M\geq N+1$ orbitals  [^2] Strictly speaking, $M=N$ would suffice but this case is often degenerate.. For strongly interacting systems, a larger number of orbitals might be required to capture many-body correlation effects. In this work, we have verified that including orbitals beyond $M=N+4$ leads to a negligible population for the time scales probed by our dynamics sup .
@@
Results for noninteracting fermions — We begin by describing the dynamics for noninteracting particles. Fig. 1(a) shows the noninteracting phase diagrams for $N=4$ fermions with $M=8$ orbitals, constructed by calculating the imbalance and expansion observables as the depth of the primary and detuning lattices are varied. We use parameter ranges compatible with experimental regimes, i.e. $V_{p}\in[3E_{r},8E_{r}]$ and $V_{d}\in[0,1.2E_{r}]$. The imbalance is calculated by averaging over the last 50 time steps at the end of each time evolution. The edge density is calculated at a target time $\bar{t}$, well before the particles can reach the boundaries of the system. Typical target times are $\bar{t}=0.45$ ms to $\bar{t}=1.78$ ms  [^3] This measurement timescale is slightly different than the experimental one, where expansion measurements are calculated at longer times. This is necessary to avoid simulating very large spatial grids that are computationally unattainable. Nevertheless, we have verified that the value of the target time $\bar{t}$ only slightly impacts the precise location of the phase boundary between the intermediate and localized phase, and that the qualitative behavior is the same for different values of $\bar{t}$.. From Fig. 1(a), we can clearly distinguish the three expected phases, which have the same structure observed in prior studies Li et al. (2017); Lueschen et al. (2018). At low values of $V_{d}$, the system is always in the extended phase (purple region). Increasing the value of $V_{d}$ at low to moderate values of $V_{p}$ pushes the system into an intermediate region where extended and localized states coexist (magenta region). Further increasing $V_{d}$ or $V_{p}$ eventually localizes all the states (yellow region).
@@
Figure 1: Behavior of the noninteracting system: (a) Phase diagram for $N=4$ fermions with $M=8$ orbitals in a 1D quasiperiodic potential. For the localized phase, the time-averaged imbalance $\mathcal{I}\right>$ is plotted in a continuous color scheme. (b)-(d) 2-RDM at $t=0.89$ ms for (a) localized phase ($V_{p}=8.0E_{r}$, $V_{d}=1.2E_{r}$), (b) intermediate phase ($V_{p}=6.0E_{r}$, $V_{d}=0.4E_{r}$), and (c) extended phase ($V_{p}=3.0E_{r}$, $V_{d}=0E_{r}$).
The three phase of the quasicrystal system can also be distinguished by the dynamics of their many-body correlations encoded in the 2-RDM, shown for selected times in Fig. 1(b)-(d) (see supplementary materials for more time snapshots sup ). The 2-RDM in the localized phase [Fig. 1(b)] does not change at all with time and merely shows the expected exchange hole at $x=x^{\prime}$ with strong many-body correlation between the sites of the initial superlattice. On the contrary, the extended phase [Fig. 1(d)] exhibits a rapid spread of correlations at very short times already. Between the two extremes of localization and expansion lies the intermediate phase [Fig. 1(c)]. The correlation pattern here reveals an almost immediate exchange correlation between neighboring sites within the populated area of the initial superlattice, and a small spread to sites just outside of it. However, the intensity of the 2-RDM is not uniform: correlations between sites in the initial superlattice structure still dominate throughout the time evolution even for longer times sup . These results indicate that the intermediate phase is characterized by two types of correlations showcasing the coexistence of localized and extended states. Localized states mediate static, short-range correlations between the initially populated sites, while extended states mediate long-range correlations spanning many sites in the quasiperiodic lattice.
@@
Fig. 2 shows how the noninteracting phase diagram changes when interactions are progressively increased [panels (a) to (c)]. At small interaction values, the intermediate phase only grows slightly while the imbalance in the localized phase decreases. As the interactions become stronger, though, the phase diagram changes drastically and in a non-monotonic fashion. For $W=0.2E_{r}$, the intermediate phase grows sizeably at the expense of most of the localized phase and part of the extended phase. Therefore, DDI stabilize the coexistence of extended and localized states in the intermediate phase. At even stronger interactions, $W=2.0E_{r}$, the intermediate phase dominates the probed parameter region. Note that in Fig. 2(c) $V_{p}$ is plotted to larger values up to $12E_{r}$. The purely extended region is instead reduced to a tiny pool around $V_{p}=3E_{r}$, $V_{d}=0E_{r}$. Most strikingly, however, a new localized region emerges at low values of $V_{d}$. Note that this region was initially part of the extended phase in the noninteracting case. The localized phases at low and high $V_{d}$ are divided by a strip of intermediate phase which persists for very large values of $V_{p}$ (as large as $V_{p}=16.0E_{r}$ – not shown here), reminiscent of the Arnold tongues found in parametrically driven systems McLachlan (1951); Magnus and Winkler (1966); Molignini et al. (2018). In this lobe, the particles undergo density oscillations around their initial pinned configuration, indicating a potential resonance between DDI and quasiperiodic potential [see Fig. 2(f)]. While both regions above and below the intermediate phase lobe show strong localization, their behavior is different. Above the lobe (high $V_{d}$, Fig. 2(d)), the particles do not move at all from their initial configuration. Below the lobe (low $V_{d}$, Fig. 2(e)), the particle density periodically leaks back and forth from the initial boundaries while retaining a strong localization inside them, indicating a dynamically self-bound state (DSBS). Such states were recently discovered in bosonic lattice models with attractive DDI Barbiero et al. (2015); Li et al. (2020, 2021a, 2021b); Korbmacher et al. (2023). Here, we reveal that DSBS can also exist for fermionic gases with repulsive DDI and away from the lattice limit $V_{d}\to\infty$.
@@
A more systematic calculation of the density dynamics, shown in Fig. 3, allows us to better discuss the behavior in each phase. The dynamics in the extended phase, localized phase, and intermediate phase is pictured respectively in the left, middle, and right columns. The rows show the dynamics for increasing interaction strength. In the extended phase [Fig. 3(a), (d), (g)], the DDI have initially little effect on the density, which maintains its linear spread in time consistent with the Lieb-Robinson bounds Tran et al. (2021). At strong interactions, however, the long-range repulsive tail in the DDI hinders the spread of correlations Bera et al. (2019), resulting in both a delayed expansion and a more chaotic density spread [Fig. 3(g)]. In the localized phase [Fig. 3(b), (e), (h))] very small oscillations can be seen at one edge of the density profile for small and intermediate interaction strengths. At strong interactions, however, this behavior is suppressed. This stabilization of localization is also observed in the intermediate phase [Fig. 3(c), (f), (i)]. At low DDI, the density in the intermediate phase maintains a strong imbalance towards the initial superlattice configuration (due to localized states) while expanding towards outer lattice sites (due to extended states). However, stronger DDI increase the imbalance while simultaneously reducing the expansion. This behavior is again consistent with a less effective spread of correlations in dipolar systems due to the interference between the repulsive tail in the interactions. Furthermore, it is an indication that DDI can be used to control the value of the SPME similarly to what was shown for contact interactions Wang et al. (2022).
@@
The increase in localization due to DDI interference is also observed in the 2-RDM. This is shown in Fig. 4 for the three different phases and at three increasing values of $W$. The behavior for weak DDI [$W=0.02E_{r}$, Fig. 4(a)-(c)] is similar to what already observed in the noninteracting case. However, when DDI are increased [$W=0.2E_{r}$ in Fig. 4(d)-(f) and $W=2.0E_{r}$ in Fig. 4(g)-(i)], the correlation spread becomes progressively stiffer for all phases. This leads to a complete freeze of correlations in the localized phase [Fig. 4(g)], while in the intermediate phase correlations are able to spread only partially  [^4] We attribute this unidirectional spread to pinning by the given choice of the quasiperiodic potential. Thus, the location of the pinning might change with a different choice of the wave vector ratio and detuning phase.. Even for the extended phase, the pinning effect is visible, with correlations spreading more easily towards the right than the left. This observation agrees with the uneven density spread observed in Fig. 3(g).